OpenStudy (anonymous):

check whether 7+3x is a factor of 3x(cube)+7x

OpenStudy (anonymous):

7+3x=0 3x=0-7 3x=-7 x=-7/3 p(-7/3)=3*(-7/3)(cube)+7*-7/3 =3*-343/27-49/3 =-343/9-49/3 -490/9

jhonyy9 (jhonyy9):

3x^3 +7x = x(3x^2 +7) - so not is a factor

OpenStudy (anonymous):

oh its not a factor